# On the non-Markovian quantum control dynamics

### v3

@@ VI Conclusion
In this paper, we study quantum control dynamics within a cavity-QED system, focusing on the non-Markovian interactions between the quantum system and its environment. The evolution of parameters representing the atom’s non-Markovian decay into the environment is described by nonlinear dynamics. The transition to Markovian interactions between the quantum system and the environment is explained by the stability of these nonlinear processes. Consequently, the dynamics of the multi-level system in a non-Markovian environment can be modeled by linear time-varying equations. Manipulation of atomic states and photons in the cavity is then achieved using open-loop and closed-loop control methods that utilize quantum measurement feedback. This approach can be extended to scenarios involving multiple coupled cavities described by high-dimensional linear time-varying equations, where feedback control can further influence the dynamics between stable and unstable subspaces.
@@ Appendix A Derivation of the non-Markovian parameter dynamics
$H_{I}=\sum_{n=1}^{N-1}\sum_{\omega}\chi_{\omega}^{(n)}\left(|n+1\rangle\langle n|b_{\omega}+|n\rangle\langle n+1|b_{\omega}^{{\dagger}}\right),$
with the detailed meaning introduced after Eq. (2) in the main text. Here we take a representative simplified case that the quantum system is coupled to the environment via operators $L_{n}$. The evolution of quantum states is influenced by its stochastic interaction with the environment as [27, 26, 41, 42]
@@
where $z_{s}$ is a complex-valued Wiener process at time $s$, $f_{n}(t,s)$ is a two-time function to be determined, related to the two-time points $t$ and $s$ [27].
Combined with Eq. (3), the differential of Eq. (76) upon the time $t$ reads
@@
$=-iH\frac{\delta|\psi\rangle}{\delta z_{s}}+if_{n}(t,s)L_{n}H|\psi\rangle+L_{n}\frac{\delta|\psi\rangle}{\delta z_{s}}z_{t}$
$~{}~{}~{}-\frac{\delta}{\delta z_{s}}L_{n}^{{\dagger}}\int_{0}^{t}\alpha(t,s)\frac{\delta|\psi\rangle}{\delta z_{s}}\mathrm{d}s-f_{n}(t,s)L_{n}L_{n}|\psi\rangle z_{t}$
@@
$=i\left[L_{n},H\right]f_{n}(t,s)|\psi\rangle$
$~{}~{}~{}+F_{n}(t)\left(L_{n}L_{n}^{{\dagger}}-L_{n}^{{\dagger}}L_{n}\right)f_{n}(t,s)L_{n}|\psi\rangle,$
@@
Then combined with Eq. (3), for a specific $L_{n}$, Eq. (78) can be rewritten as
$\frac{\partial f_{n}(t,s)}{\partial t}L_{n}|\psi\rangle$ (79)
$=$ $i\omega_{n}f_{n}(t,s)L_{n}|\psi\rangle+F_{n}(t)\kappa_{n}f_{n}(t,s)L_{n}|\psi\rangle,$
further combined with Eq. (7), we have
$\dot{F}_{n}(t)=\frac{\gamma\chi_{n}}{2}-(\gamma+i\Omega)F_{n}(t)$ (80)
$~{}~{}~{}~{}+\int_{0}^{t}\alpha_{n}(t,s)\left[i\omega_{n}f_{n}(t,s)+F_{n}(t)\kappa_{n}f_{n}(t,s)\right]\mathrm{d}s$
@@
where $f_{n}(t,t)=\chi_{n}$ is a constant, $F(0)=0$, and the analytic solutions of generalized $F_{n}(t)$ is given in Ref. [27].
@@ Appendix B Derivation of the non-Markovian master equation
The non-Markovian master equation has been introduced in Refs. [27, 29]. In this Appendix, we briefly introduce the derivation of the non-Markovian control equation (6) in the main text, which is the control equation we focus on in this paper.
The evolution of $f_{n}(t,s)$ has been clarified in Appendix A. Combining Eq. (4) with Eq. (76) after the normalization, we have the following equation for the state vector,
$\frac{\mathrm{d}}{\mathrm{d}t}|\psi(t)\rangle=-iH|\psi(t)\rangle+\sum_{n=1}^{N-1}\left(L_{n}-\left\langle L_{n}\right\rangle\right)|\psi(t)\rangle z_{t}$ (81)
$-\sum_{n=1}^{N-1}\left(L_{n}^{{\dagger}}-\left\langle L_{n}^{{\dagger}}\right\rangle\right)\int_{0}^{t}\alpha(t,s)f_{n}(t,s)L_{n}|\psi(t)\rangle\mathrm{d}s.$
@@
$\dot{\rho}=-i[H,\rho]+\sum_{n=1}^{N-1}\left\{\int_{0}^{t}\mathrm{d}sE_{z}\left[z_{t}z_{s}^{*}\right]f_{n}^{*}(t,s)+\right.$ (82)
$\left.\int_{0}^{t}\mathrm{d}sE_{z}\left[z_{t}^{*}z_{s}\right]f_{n}(t,s)L_{n}\rho L_{n}^{{\dagger}}-\int_{0}^{t}\alpha(t,s)f_{n}(t,s)\mathrm{d}s\right.$
$\left.L_{n}^{{\dagger}}L_{n}\rho-\rho L_{n}^{{\dagger}}L_{n}\int_{0}^{t}\alpha^{*}(t,s)f_{n}^{*}(t,s)\mathrm{d}s\right\},$
and can be further written as Eq. (6) in the main text, which agrees with the format in Ref. [29].
@@ Appendix C Some results on nonlinear and linear dynamics
Suppose for each $u\geq 0$, there exists a positive Lyapunov function $V_{u}$ for the system with the input $u$ and
$\dot{V}_{u}(t,X_{f})\leq 0.$
When the input satisfies $\|u\|\leq a_{u}$, then the system is BIBO stable.
@@
where f is continuously differentiable with $\textbf{f}(0,t)=0$ and $\textbf{h}(\textbf{X}(t),t)$ is a persistent perturbation: if the system is uniformly and asymptotically stable about its equilibrium $\textbf{X}^{*}=0$ when $\textbf{h}(\textbf{X}(t),t)=0$, and there are two positive constants $\tilde{\delta}_{1}$ and $\tilde{\delta}_{2}$ such that $\|\textbf{h}(\textbf{X}(t),t)\|<\tilde{\delta}_{1}$ for $t\in[0,\infty]$ and $\|\textbf{h}(\textbf{X}(0),0)\|<\tilde{\delta}_{2}$, then the persistently perturbed system remains to be stable in the sense of Lyapunov.
@@
A linear time varying system is uniformly exponential stable (UES) if and only if there exist positive constants $K$ and $\tilde{\alpha}$ such that
$\|\Phi\left(t,\tau\right)\|\leq Ke^{-\tilde{\alpha}\left(t-\tau\right)},$
for $t_{0}\leq\tau<t<\infty$.
@@ Acknowledgements
This work is supported by the ANR project “Estimation et controle des systèmes quantiques ouverts” Q-COAST Project ANR- 19-CE48-0003 and the ANR project IGNITION ANR-21-CE47-0015.

### v4

@@ VI Conclusion
In this paper, we study quantum control dynamics based on a cavity-QED system, focusing on the non-Markovian interactions between the quantum system and its environment. The evolution of parameters representing the atom’s non-Markovian decay into the environment is described by nonlinear dynamics. The transition to Markovian interactions between the quantum system and the environment is explained by the stability of these nonlinear processes. Consequently, the dynamics of the multi-level system in a non-Markovian environment can be modeled by linear time-varying equations. Manipulation of atomic states and photons in the cavity is then achieved using open-loop and closed-loop control methods that utilize quantum measurement feedback. This approach can be extended to scenarios involving multiple coupled cavities described by high-dimensional linear time-varying equations, where feedback control can further influence the dynamics between stable and unstable subspaces.
@@ Appendix A Derivation of the non-Markovian parameter dynamics
$H_{\rm AE}=\sum_{n=1}^{N-1}\sum_{\omega}\chi_{\omega}^{(n)}\left(|n+1\rangle\langle n|b_{\omega}+|n\rangle\langle n+1|b_{\omega}^{{\dagger}}\right),$
and the detailed meaning is introduced after Eq. (2) in the main text. Here we take a representative simplified case that the quantum system is coupled to the environment via operators $L_{n}$. The evolution of quantum states is influenced by its stochastic interaction with the environment as [29, 28, 43, 44]
@@
where $z_{s}$ is a complex-valued Wiener process at time $s$, $f_{n}(t,s)$ is a two-time function to be determined, related to the two-time points $t$ and $s$ [29]. Combined with Eq. (3), the differential of Eq. (74) upon the time $t$ reads
@@
$=-iH\frac{\delta|\psi\rangle}{\delta z_{s}}+if_{n}(t,s)L_{n}H|\psi\rangle+L_{n}\frac{\delta|\psi\rangle}{\delta z_{s}}z_{t}-\frac{\delta}{\delta z_{s}}L_{n}^{{\dagger}}\int_{0}^{t}\alpha(t,s)\frac{\delta|\psi\rangle}{\delta z_{s}}\mathrm{d}s-f_{n}(t,s)L_{n}L_{n}|\psi\rangle z_{t}$
@@
$=i\left[L_{n},H\right]f_{n}(t,s)|\psi\rangle+F_{n}(t)\left(L_{n}L_{n}^{{\dagger}}-L_{n}^{{\dagger}}L_{n}\right)f_{n}(t,s)L_{n}|\psi\rangle,$
@@
Then combined with Eq. (3), for a specific $L_{n}$, Eq. (76) can be rewritten as $\partial_{t}f_{n}(t,s)L_{n}|\psi\rangle=i\omega_{n}f_{n}(t,s)L_{n}|\psi\rangle+F_{n}(t)\kappa_{n}f_{n}(t,s)L_{n}|\psi\rangle$, further combined with Eq. (7), we have
$\dot{F}_{n}(t)$ $=\frac{\gamma\chi_{n}}{2}-(\gamma+i\Omega)F_{n}(t)+\int_{0}^{t}\alpha(t,s)\left[i\omega_{n}f_{n}(t,s)+F_{n}(t)\kappa_{n}f_{n}(t,s)\right]\mathrm{d}s$ (77)
@@
where $f_{n}(t,t)=\chi_{n}$ [29].
@@ Appendix B Derivation of the non-Markovian master equation
The non-Markovian master equation has been introduced in Refs. [29, 31]. In this Appendix, we briefly introduce the derivation of the non-Markovian control equation (6) in the main text. Based on $f_{n}(t,s)$ clarified in Appendix A, combine Eq. (4) with Eq. (74) after the normalization, we have the following equation for the state vector,
$\frac{\mathrm{d}}{\mathrm{d}t}|\psi(t)\rangle=-iH|\psi(t)\rangle+\sum_{n=1}^{N-1}\left(L_{n}-\left\langle L_{n}\right\rangle\right)|\psi(t)\rangle z_{t}-\sum_{n=1}^{N-1}\left(L_{n}^{{\dagger}}-\left\langle L_{n}^{{\dagger}}\right\rangle\right)\int_{0}^{t}\alpha(t,s)f_{n}(t,s)L_{n}|\psi(t)\rangle\mathrm{d}s.$ (78)
@@
$\dot{\rho}=$ $-i[H,\rho]+\sum_{n=1}^{N-1}\left\{\int_{0}^{t}\mathrm{d}sE_{z}\left[z_{t}z_{s}^{*}\right]f_{n}^{*}(t,s)+\int_{0}^{t}\mathrm{d}sE_{z}\left[z_{t}^{*}z_{s}\right]f_{n}(t,s)L_{n}\rho L_{n}^{{\dagger}}\right$. (79)
$\left.-\int_{0}^{t}\alpha(t,s)f_{n}(t,s)\mathrm{d}sL_{n}^{{\dagger}}L_{n}\rho-\rho L_{n}^{{\dagger}}L_{n}\int_{0}^{t}\alpha^{*}(t,s)f_{n}^{*}(t,s)\mathrm{d}s\right\},$
and can be further written as Eq. (6) in the main text [31].
@@ Appendix C Some results on nonlinear and linear dynamics
Suppose for each $u\geq 0$, there exists a positive Lyapunov function $V_{u}$ for the system with the input $u$ and $\dot{V}_{u}(t,X_{f})\leq 0$. When the input satisfies $\|u\|\leq a_{u}$, the system is BIBO stable.
@@
where f is continuously differentiable with $\textbf{f}(0,t)=0$ and $\textbf{h}(\textbf{X}(t),t)$ is a persistent perturbation: if the system is uniformly and asymptotically stable about its equilibrium $\textbf{X}^{*}=0$ when $\textbf{h}(\textbf{X}(t),t)=0$, and there are two positive constants $\tilde{\delta}_{1}$ and $\tilde{\delta}_{2}$ such that $\|\textbf{h}(\textbf{X}(t),t)\|<\tilde{\delta}_{1}$ for $t\in[0,\infty)$ and $\|\textbf{h}(\textbf{X}(0),0)\|<\tilde{\delta}_{2}$, then the persistently perturbed system remains to be stable in the sense of Lyapunov.
@@
A linear time varying system is uniformly exponential stable (UES) if and only if there exist positive constants $K$ and $\tilde{\alpha}$ such that $\|\Phi\left(t,\tau\right)\|\leq Ke^{-\tilde{\alpha}\left(t-\tau\right)}$, for $t_{0}\leq\tau<t<\infty$.
@@ Acknowledgements
This work is supported by the ANR project “Estimation et controle des systèmes quantiques ouverts” Q-COAST Project ANR- 19-CE48-0003 and the ANR project IGNITION ANR-21-CE47-0015, Quantum Science and Technology-National Science and Technology Major Project 2023ZD0300600, Guangdong Provincial Quantum Science Strategic Initiative No. GDZX2303007, and Hong Kong Research Grant Council (RGC) under Grant No. 15213924.